• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 20:38:47 on June 3,2002

    Entry number 85169

    keyword=Target reboot

    Target IOC had to be rebooted, while beam was on at 100uA. The beam did not trip during the reboot. Everything's OK.